OpenStudy (anonymous):

In an experiment a petri dish with a colony of bacteria is exposed to cold temperatures and then warmed again. Find a quadratic model for the data in the table. Show work

OpenStudy (anonymous):

Getting the table, one moment

OpenStudy (anonymous):

Time (hours) 0 1 2 3 4 5 6 Population (1000s) 5.1 3.03 1.72 1.17 1.38 2.35 4.08

OpenStudy (anonymous):

how are you supposed to do this (just asking because you have to "show work")

OpenStudy (anonymous):

I have no clue lol I typed everything the question asked and the table

OpenStudy (decentnabeel):

The general formula for quadratic eqn is y = ax^2 + bx + c take any 3 points from the table (0, 5.1) (1, 3.03) (5, 2.35) plug these points into the above equation and you will get 3 different equations use substitution or any method that you know to solve these 3 eqns with 3 unknowns

OpenStudy (decentnabeel):

a point is written in the form (x,y) so plug 0 for x and 5.1 for y to get 1st eqn. follow same method to get other 2 eqns

OpenStudy (anonymous):

@DecentNabeel i am not sure that would work, since it is not exact

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

So this part is correct? The general formula for quadratic eqn is y = ax^2 + bx + c take any 3 points from the table (0, 5.1) (1, 3.03) (5, 2.35) plug these points into the above equation and you will get 3 different equations use substitution or any method that you know to solve these 3 eqns with 3 unknowns

OpenStudy (decentnabeel):

@satellite73 many way for solve i solve another but this is true

OpenStudy (anonymous):

if it is exact you can do it using only three points \[(0,5.1),(1,3.03),(2,2.35)\] or any other 3

OpenStudy (decentnabeel):

yes is correct

OpenStudy (anonymous):

also you do not have to solve a system of three equations, only two

OpenStudy (anonymous):

you want \[ax^2+bx+c\] but sine \((0,5.1\) is on the graph, you have \(c=5.1\) right away, only need \(a\) and \(b\)

OpenStudy (anonymous):

So lets write step one so i can start writing it so i can show our work lol

OpenStudy (anonymous):

the part you wrote above is correct

OpenStudy (decentnabeel):

5.1 = a*0^2 + b*0 + c 3.03 = a*1^2 + b*1 + c 2.35 = a*5^2 + b*5 + c solve of a, b and c

OpenStudy (anonymous):

what @DecentNabeel said solve that system but really it is not as hard as it looks since the first one tells you \(c=5.1\)

OpenStudy (anonymous):

I dont know how to plug all that in

OpenStudy (anonymous):

oops second one is \[3.03=a+b=5.1\] so \[a+b=-2.07\]

OpenStudy (anonymous):

do you see how @DecentNabeel got this one \[5.1 = a*0^2 + b*0 + c\]

OpenStudy (anonymous):

if the answer is "no" that is fine, just say so

OpenStudy (decentnabeel):

yes c=5.1 @satellite73

OpenStudy (anonymous):

no i dont get it

OpenStudy (anonymous):

ok do you get that you are trying to find \(a,b\) and \(c\) to make up \[y=ax^2+bx+c\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok and from the table you see that \((0,5.1)\) is on the graph right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

that means if \(x=0\) then \(y=5.1\) right ?

OpenStudy (anonymous):

correct

OpenStudy (decentnabeel):

What @LynFran ????

OpenStudy (anonymous):

@satellite73 are you back? The serve kicked me off

OpenStudy (anonymous):

yea me too

OpenStudy (decentnabeel):

@EllenJaz17 what are you said @LynFran

OpenStudy (anonymous):

they interrupted me learning lol

OpenStudy (anonymous):

lol

OpenStudy (decentnabeel):

lol

OpenStudy (anonymous):

ok so to resume our story, were were here: we want \(a,b,c\) so make \(y=ax^2+bx+c\) and we know if \(x=0\) then \(y=5.1\) plug them in and get \[5.1=a\times 0^2+b\times 0+c\]

OpenStudy (anonymous):

is that ok?

OpenStudy (anonymous):

yes i get that one because its easy because x is 0 lol

OpenStudy (anonymous):

right so that tells us right away that \(c=5.1\) yes?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ok now \(a\) and \(b\) to go

OpenStudy (anonymous):

@LynFran send him a message, im trying to get hw help, not get interrupted

OpenStudy (anonymous):

ok im ready

OpenStudy (anonymous):

the next point we have is \((1,3.03)\) right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now don't forget, we already know \(c=5.1\) so we have \[y=ax^2+bx+5.1\] put \(x=1\) and \(y=3.03\) to get \[3.03=a\times 1^2+b\times 1+5.1\]

OpenStudy (anonymous):

this also cleans up nicely, since \(1^2=1\) we have \[3.03=a+b+5.1\]

OpenStudy (anonymous):

good so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now to make it neater, subtract \(5.1\) from both sides that gives \[-2.07=a+b\] or \[a+b=-2.07\] and we put that aside for the moment

OpenStudy (anonymous):

how we doing so far?

OpenStudy (anonymous):

good

OpenStudy (anonymous):



OpenStudy (anonymous):

ok now one more equation to go you can pick any point you like but i will pick \((2,1.72)\)

OpenStudy (anonymous):

Sounds good

OpenStudy (decentnabeel):

good @satellite73

OpenStudy (anonymous):

again we have \[y=ax^2+bx+5.1\] put \(x=2,y=1.72\) and get \[1.72=a\times 2^2+b\times 2+5.1\]

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we clean it up actually you can what do you get in terms of \(a,b\) and a number ?

OpenStudy (anonymous):

\[-2.07^{2}+1.72+5.1\] ? Im sure that's wrong

OpenStudy (anonymous):

quite wrong, it is much easier than that

OpenStudy (anonymous):

haha oh jeez. I needed this class apparently very much lol and your help

OpenStudy (anonymous):

lets go back to \[1.72=a\times 2^2+b\times 2+5.1\]

OpenStudy (anonymous):

your equation will have an \(a\) and a \(b\) in it all you have to do is compute \(2^2\)

OpenStudy (anonymous):

So a=4?

OpenStudy (anonymous):

no just \(a\times 4\) or \(4a\)

OpenStudy (anonymous):

lets back up a second

OpenStudy (anonymous):

remember that when we replace \(x\) by \(1\) and \(y\) by \(3.03\) we ended up with an equation with both an \(a\) and a \(b\) in it it was \[a+b=-2.07\]

OpenStudy (anonymous):

we are trying to accomplish the same thing here with \(x=2\) and \(y=1.72\) first step is to write \[1.72=a\times 2^2+b\times 2+5.1\]

OpenStudy (anonymous):

so don't do too much work just see what the numbers give you your equation should still have an \(a\) and a \(b\) in it

OpenStudy (anonymous):

if you are still confused, say so and i will spell it out

OpenStudy (anonymous):

Yeah im confused

OpenStudy (anonymous):

thought so i am just going to rewrite \[1.72=a\times 2^2+b\times 2+5.1\]

OpenStudy (anonymous):

since \(2^2-4\) this is the same as \[1.72=a\times 4+b\times 2+5.1\] clear?

OpenStudy (anonymous):

yes